Determine the Q point of the circuit shown in fig. Also, draw the D.C load line. Given β=158
& RC = 1058 Ω.
1
Expert's answer
2020-10-05T09:50:40-0400


According to KVL & KCL:


"I_BR_B+0.7+I_ER_E=10,\\
I_E=I_C+I_B."


Since this is an amplifier:


"V_B>V_E,\\
V_B<V_C.\\
I_C=\beta I_B\rightarrow I_E=(1+\beta)I_B,\\
I_BR_B+(1+\beta)I_BR_E=10-0.7,\\
I_B=\textbf{0.01171}\text{ mA},\\
I_E=(1+\beta)I_B=\textbf{1.862}\text{ mA},\\
I_C=I_E-I_B=\textbf{1.850}\text{ mA}."

Voltages:


"V_{CC}=I_CR_C+V_C,\\
10=I_CR_C+V_C,\\
V_C=10-I_CR_C=\textbf{8.043}\text{ V}.\\
V_E=I_ER_E-10=\textbf{-1.249}\text{ V}.\\
V_B=V_{BE}+V_E=0.7-1.249=\textbf{-0.549}\text{ V}."

Q point, therefore, is:


"(V_{C},I_C)=(8.043\text{ V},1.850\text{ mA})"

For DC load line the saturation current:


"I{_C(sat)}=\frac{V_{CC}-V_E}{R_C}=0.0106\text{ A}."

Slope:


"S=\frac{-0.0106}{10-(-1.249)}=-9.42\cdot10^{-4}\space\Omega^{-1}."